# Learning large softmax mixtures with warm start EM

### v1

@@ 4 Simulations
In this section we conduct numerical experiments to corroborate our theoretical findings in Section 2 and Section 3. To generate the data, we first generate $X_{1},\ldots,X_{p}$ i.i.d. from $\mathcal{N}_{L}(0,\mathbf{I}_{L})$. The mixing weights are set to ${\bm{\alpha}}^{*}=(1/K)1_{K}$. To generate the parameters ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$, we first draw a $L\times K$ matrix with entries i.i.d. from $\mathcal{N}(0,1)$. We then set its left $K$ singular vectors as ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$. Finally, we resample $Y_{1},\ldots,Y_{N}$ according to model (4).
We consider the following estimation methods: (1) MoM, the Method of Moments estimator in Section 2.3 with $n=200$ for choosing the projection direction as discussed in Remark 6; (2) EM-MoM, the EM estimator in Section 3 that uses the MoM estimator as initialization; (3) EM-rand, the EM estimator that uses random initialization; [^2] Entries of ${\bm{\alpha}}^{(0)}$ are i.i.d. drawn from Unif($0,1$) and normalized to the unit $\ell_{1}$-norm while entries of $\widehat{\bm{\theta}}_{k}^{(0)}$, $k\in[K]$, are i.i.d. from $\mathcal{N}(0,1)$ and then normalized such that its $\ell_{2}$-norm is no greater than $1$. (4) EM-rand-multi, the EM estimator that uses the initialization which maximizes the likelihood among 100 random initializations; (5) EM-oracle, the EM estimator that uses the true parameter values as the initialization. For the EM-algorithm in Section 3, we choose the step size $\eta_{k}=1$ and return the estimates after 500 iterations in order to ensure fair comparison among the above EM estimators.
@@
where $\varrho:[K]\to[K]$ is the best permutation that minimizes $\text{Err}_{{\bm{\theta}}}$. The baseline setting is chosen as $K=3$, $L=50$, $p=5000$ and $N=5000$. We vary $N\in\{500,1000,3000,5000,7000\}$, $p\in\{1000,2000,3000,5000,7000\}$ and $L\in\{10,30,50,70,100\}$ one at a time to examine their effects on the estimation errors. When $p$ is varied, we set $N=7000$. When varying $L$, we increased both $p$ and $N$ to $7000$. For each setting, we report the averaged errors over $200$ repetitions in Fig. 1 for $\text{Err}_{\bm{\theta}}$ and in Fig. 2 for $\text{Err}_{\bm{\alpha}}$.
Regarding $\text{Err}_{\bm{\theta}}$, all methods perform better as $N$ increases and $L$ decreases. For EM-oracle, since it has no algorithmic error, our Theorem 1 shows that its estimation error is purely the statistical error which is of order $\sqrt{L\log(N)/N}$. Between MoM and EM-MoM, MoM is a safer choice for small $N$ while EM-MoM has better performance in almost all settings as long as $N$ is moderate ($N\geq 2000$). Once $N\geq p$, further increasing $N$ does not improve the performance of MoM. When $p$ increases, the performance of MoM improves whereas EM-MoM remains unchanged. We also note that the figures in which we vary $N$ and $p$ suggest the rate for MoM is slower than the parametric rate, confirming the observation made in Remark 5 above. We observe similar findings for estimating the mixing weights ${\bm{\alpha}}^{*}$ except that $\text{Err}_{\bm{\alpha}}$ of MoM is not affected by increasing $L$. The large fluctuation of $\text{Err}_{{\bm{\alpha}}}$ for MoM may be due to the sensitivity of the method to the selection of the random projection in Section 2.5.
We conduct a separate simulation study below to verify that MoM can indeed enjoy a parametric rate, by taking $K=2$ and ensuring that the atoms have the theoretically prescribed separation. In addition to setting $K=2$, we let $L=50$ and vary $p=N$ within $\{1000,3000,5000,7000,9000,12000,15000\}$. Fig. 3 depicts the estimation errors of MoM, EM-MoM and EM-oracle. We observe the same phenomenon as above except that $\text{Err}_{{\bm{\theta}}}$ of MoM decays in the faster parametric rate as $N$ and $p$ increase. The large fluctuation of $\text{Err}_{{\bm{\alpha}}}$ for MoM can be explained by the sensitivity of the method to the selection of the random projection in Section 2.5.
These findings are all aligned with our theory in Section 2 and Section 3. Among the variants of EM estimators, EM-MoM and EM-rand-multi have overall comparable performance, and yield more satisfactory results than EM-rand. One drawback of EM-rand-multi is its higher computational cost due to sampling multiple initializations and evaluating their likelihood (the computation complexity scales linearly in the number of initializations multiplied by the ambient dimension $p$).
Overall, in the large $N$, $p$ regime, we recommend the usage of EM-MOM as the estimator with the best computational/statistical trade-offs, backed up by our theoretical results.

### v2

@@ 4 Simulations
In this section we conduct numerical experiments to corroborate our theoretical findings in Sections 3 and 2. In Section 4.1 we first examine how the performance of the EM and MoM estimators depends on $N$, $p$, $L$ and $K$. In Section 4.2 we demonstrate the benefit of using the dimension reduction technique from Section 3.4.2 to initialize the EM algorithm.
To generate the data, we first generate $X_{1},\ldots,X_{p}$ i.i.d. from $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$. The mixing weights are set to ${\bm{\alpha}}^{*}=(1/K)1_{K}$ for any given integer $K$. To generate the parameters ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$, we first draw a $L\times K$ matrix with entries i.i.d. from $\mathcal{N}(0,1)$. We then set its $K$ left-singular vectors as ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$. Finally, we resample $Y_{1},\ldots,Y_{N}$ according to model (2).
We consider the following estimation methods:
  • (1)
    MoM, the Method of Moments estimator in Section 3.3 with $n=200$ for choosing the projection direction as discussed in Remark 9;
  • (2)
    EM-MoM, the EM estimator in Section 2 that uses the MoM estimator as initialization;
  • (3)
[…]
  • (4)
    EM-oracle, the EM estimator that uses the true parameter values as the initialization.
For the EM algorithm in Section 2, we choose the step size $\eta_{k}=0.2$ and use the stopping rule that the relative change of the log-likelihood is smaller than $10^{-6}$.
@@
where $\varrho:[K]\to[K]$ is the best permutation that minimizes $\text{Err}_{{\bm{\theta}}}$.
4.1 Dependence of estimation error on $N,p,L$ and $K$
We vary $N\in\{2,4,6,8,10\}\times 10^{3}$, $p\in\{1,3,5,7,10\}\times 10^{3}$, $L\in\{20,40,60,80,100\}$ and $K\in\{2,4,6,8,10\}$ one at a time to examine their effects on the estimation errors. The baseline setting uses $L=50$ and $K=3$ when these parameters are not varied. When $N$ is varied, we set $p=5000$ and when $p$ is varied, we set $N=7000$. When varying either $L$ or $K$, we chose $N=10000$ and $p=7000$. For each setting, we report the averaged errors over 200200200 repetitions in Fig. 1 for $\text{Err}_{\bm{\theta}}$ (and in Fig. 3 of Section A.1 for $\text{Err}_{\bm{\alpha}}$).
Regarding $\text{Err}_{\bm{\theta}}$, all methods perform better as $N$ increases and $L$ or $K$ decreases. For EM-oracle, since it has no algorithmic error, our Theorem 2 shows that its estimation error is purely the statistical error which is of order $\sqrt{L\log(N)/N}$. The MoM estimator is outperformed by the EM estimators in all settings. Once $N\geq p$, further increasing $N$ does not improve the performance of MoM. When $p$ increases, the performance of MoM improves, whereas the EM estimators remain unchanged. We also note that the figures in which we vary $N$ and $p$ suggest the rate for MoM is slower than the parametric rate, confirming the observation made in Remark 11 above. In Section A.2, we conduct a separate simulation study below to verify that MoM can indeed enjoy a parametric rate.
Overall, for $K=3$, EM-MoM and EM-dr-rand-10 have overall comparable performance, with the former performing slightly better for large $N$. One drawback of EM-dr-rand-10 is its higher computational cost due to sampling multiple initializations and evaluating their likelihoods (the computational complexity scales linearly with the number of initializations multiplied by the ambient dimension $p$).
As $K$ increases, the performance of all methods deteriorates, with MoM and EM-MoM degrading more rapidly than the others. For $K=10$, MoM (so does EM-MoM) fails to recover all $K$ mixture components, as the root-finding step in Algorithm 1 fails in this case.
These findings are all aligned with our theory in Sections 3 and 2.
@@ 4.1 Dependence of estimation error on $N,p,L$ and $K$
4.2 Benefits of multiple random initializations with dimension reduction
[…]
Figure 2: The averaged $\text{Err}_{{\bm{\theta}}}$ in different settings
Appendix A Additional simulations
A.1 Results of estimating $\alpha^{*}$ in the setting of Section 4.1
Fig. 3 shows the errors $\text{Err}_{\bm{\alpha}}$ of all methods in Section 4.1. For estimating the mixing weights ${\bm{\alpha}}^{*}$, both MoM and EM-MoM exhibit greater fluctuations in their errors due to the method’s sensitivity to the choice of random projection in Section 3.4.1. For large $K$, the errors in estimating ${\bm{\alpha}}^{*}$ are substantially larger for these methods compared to other EM estimators, and are more sensitive than the corresponding errors in estimating ${\bm{\theta}}_{k}^{*}$. EM-dr-rand-10 and EM-oracle perform better for larger $N$ and smaller $K$, whereas their performance remains similar as $p$ and $L$ vary.
@@ Appendix A Additional simulations
A.2 Parametric rate of MoM
We conduct a separate simulation study below to verify that MoM can indeed enjoy a parametric rate, by taking $K=2$ and ensuring that the atoms have the theoretically prescribed separation. We let $L=50$ and vary $p=N\in\{1,3,5,7,9,12,15\}\times 10^{3}$. Fig. 4 depicts the estimation errors of MoM, EM-MoM and EM-oracle. We observe the same phenomenon as above except that $\text{Err}_{{\bm{\theta}}}$ of MoM decays in the faster parametric rate as $N$ and $p$ increase. The large fluctuation of $\text{Err}_{{\bm{\alpha}}}$ for MoM can be explained by the sensitivity of the method to the selection of the random projection in Section 3.4.1.